Find 2×2 matrix A that maps (1,3)^T and (1,4)^T into (-2,5)^T and (3,-1)^T, respectively

1
Expert's answer
2021-11-30T14:04:08-0500

Solution;

A linear transformation "T_A:R^2\to R^2" Is defined by;

"T_A(v)=A(v)"

Since A is the desired matrix of transformation,we have;

"A\begin{bmatrix}
 1\\
 3
\end{bmatrix}=\begin{bmatrix}
 -2 \\
 5
\end{bmatrix}" and "A\begin{bmatrix}
 1 \\
 4
\end{bmatrix}=\begin{bmatrix}
 3 \\
 -1
\end{bmatrix}"

If we put both maps into one matrix ,we have;

"A\begin{bmatrix}
 1 & 1\\
 3 & 4
\end{bmatrix}=\begin{bmatrix}
 -2 & 3 \\
 5 & -1
\end{bmatrix}"

Hence we can find A as;

"A=\begin{bmatrix}
 -2 & 3 \\
 5 & -1
\end{bmatrix}\begin{bmatrix}
 1 & 1 \\
 3 & 4
\end{bmatrix}^{-1}"


"A=\begin{bmatrix}
 -2 & 3\\
 5& -1
\end{bmatrix}\begin{bmatrix}
 4 & -1 \\
 -3 & 1
\end{bmatrix}"


"A=\begin{bmatrix}
 -17& 5\\
 23 & -6
\end{bmatrix}"